a rectangular playground at a park measures 62 feet by 48 feet. there is a 4.5 foot wide walkway that borders the perimeter of the playground. what is the area of the walkway
Answer by ikleyn(52932)

.
The area of the walkway = area of the large rectangle - area of the park. 


    The dimensions of the large rectangle are  62+2*4.5 = 71 ft  and  48+2*4.5 = 57 ft.


Therefore, the area under the question is  71*57 - 62*48 = 1071 square feet.

Solved.